# Bayesian Optimization for Non-Convex Two-Stage Stochastic Optimization Problems

### v1

@@ A.3 Uniform convergence of the posterior mean and covariance functions
In this section, we establish the uniform convergence of the posterior mean and covariance functions of $f$ and $\overline{f}$. This will be an important ingredient in proving convergence of the values of the recommendations.
@@
The proof uses the theory of martingales in Banach spaces. The Gaussian process, $f$, is defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$ and can be viewed as a function $f:\Omega\times\tilde{{\mathbb{X}}}\to\mathbb{R},\,(\omega,\tilde{{\bm{x}}})\mapsto f(\tilde{{\bm{x}}})(\omega)$. The sigma-algebra $\mathcal{F}$ contains all the events generated by the finite dimensional distributions of $f$ and, since sigma-algebras are closed under countable intersections, $\mathcal{F}$ also contains all events which can be described by the value of $f$ at countably many locations in $\tilde{{\mathbb{X}}}$. We are considering the version of $f$ with continuous sample paths and $\tilde{{\mathbb{X}}}$ is separable, and so $\mathcal{F}$ contains all events described by the sample paths of $f$. That is, we can view $f$ as a measurable function $f:\Omega\to\mathcal{C}(\tilde{{\mathbb{X}}}),\,\omega\mapsto(\tilde{{\bm{x}}}\mapsto f(\omega,\tilde{{\bm{x}}}))$ taking values in the space $\mathcal{C}(\tilde{{\mathbb{X}}})$ of continuous, real-valued functions on $\tilde{{\mathbb{X}}}$. In other words, $f$ is a random element in $\mathcal{C}(\tilde{{\mathbb{X}}})$, defined on $(\Omega,\mathcal{F},\mathbb{P})$.
@@
We now prove the analogous result of Proposition A.11 for the average performance objective, $\overline{f}$. For each $n\in\mathbb{N}_{0}\cup\{\infty\}$, ${\bm{x}},{\bm{x}}^{\prime}\in{\mathbb{X}}$ and ${\bm{g}},{\bm{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}$, define the posterior mean and covariance functions of $\overline{f}$ as
$\overline{\mu}^{n}({\bm{x}},{\bm{g}})$ $=\mathbb{E}[\overline{f}({\bm{x}},{\bm{g}})\,|\,\mathcal{F}^{n}],$ (30a)
$\overline{k}^{n}({\bm{x}},{\bm{g}};\,{\bm{x}}^{\prime},{\bm{g}}^{\prime})$ $=\mathrm{Cov}[\overline{f}({\bm{x}},{\bm{g}}),\overline{f}({\bm{x}}^{\prime},{\bm{g}}^{\prime})\,|\,\mathcal{F}^{n}].$ (30b)
Note that we do not include a statement of continuity of $\overline{\mu}^{\infty}$ and $\overline{k}^{\infty}$ since we have not defined a topology on the set ${\mathbb{X}}\times\{{\mathbb{U}}\to{\mathbb{Y}}\}$.
@@
For any choice of design points $\tilde{{\mathbf{x}}}^{1},\tilde{{\mathbf{x}}}^{2},\dots$, the sequences of stochastic processes $\overline{\mu}^{n}\to\overline{\mu}^{\infty}$ and $\overline{k}^{n}\to\overline{k}^{\infty}$ converge uniformly as $n\to\infty$, both almost surely and in $L^{p}$ for all $1\leq p<\infty$.
@@
$\sup_{\begin{subarray}{c}{\mathbf{x}}\in{\mathbb{X}}\\ {\mathbf{g}}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\bigl{|}\overline{\mu}^{n}({\mathbf{x}},{\mathbf{g}})-\overline{\mu}^{\infty}({\mathbf{x}},{\mathbf{g}})\bigr{|}\leq\sup_{\begin{subarray}{c}{\mathbf{x}}\in{\mathbb{X}}\\ {\mathbf{y}}\in{\mathbb{Y}}\\ {\mathbf{u}}\in{\mathbb{U}}\end{subarray}}\bigl{|}\mu^{n}({\mathbf{x}},{\mathbf{y}},{\mathbf{u}})-\mu^{\infty}({\mathbf{x}},{\mathbf{y}},{\mathbf{u}})\bigr{|}\to 0$
@@
$\sup_{\begin{subarray}{c}{\mathbf{x}},{\mathbf{x}}^{\prime}\in{\mathbb{X}}\\ {\mathbf{g}},{\mathbf{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\bigl{|}\overline{k}^{n}({\mathbf{x}},{\mathbf{g}};\,{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})-\overline{k}^{\infty}({\mathbf{x}},{\mathbf{g}};\,{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\bigr{|}\leq\sup_{\tilde{{\mathbf{x}}},\tilde{{\mathbf{x}}}^{\prime}\in\tilde{{\mathbb{X}}}}\bigl{|}k^{n}(\tilde{{\mathbf{x}}},\tilde{{\mathbf{x}}}^{\prime})-k^{\infty}(\tilde{{\mathbf{x}}},\tilde{{\mathbf{x}}}^{\prime})\bigr{|}\to 0$
@@
All moments of the supremum of $|\overline{f}|$ are finite. That is,
 $\forall\,1\leq p<\infty\qquad\mathbb{E}\Biggl{[}\biggl{(}\sup_{\begin{subarray}{c}{\mathbf{x}}\in{\mathbb{X}}\\ {\mathbf{g}}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}|\overline{f}({\mathbf{x}},{\mathbf{g}})|\biggr{)}^{p}\Biggr{]}<\infty.$
@@
$\mathbb{E}\Biggl{[}\biggl{(}\sup_{\begin{subarray}{c}{\mathbf{x}}\in{\mathbb{X}}\\ {\mathbf{g}}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}|\overline{f}({\mathbf{x}},{\mathbf{g}})|\biggr{)}^{p}\Biggr{]}=\mathbb{E}\Biggl{[}\biggl{(}\sup_{\begin{subarray}{c}{\mathbf{x}}\in{\mathbb{X}}\\ {\mathbf{g}}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\Bigl{|}\mathbb{E}_{\mathbf{u}}\bigl{[}f({\mathbf{x}},{\mathbf{g}}({\mathbf{u}}),{\mathbf{u}})\bigr{]}\Bigr{|}\biggr{)}^{p}\Biggr{]}\leq\mathbb{E}\biggl{[}\sup_{{\mathbf{x}},{\mathbf{y}},{\mathbf{u}}}|f({\mathbf{x}},{\mathbf{y}},{\mathbf{u}})|^{p}\biggr{]}<\infty.$
@@ A.4 Convergence to optimal recommendations
The main result in this section is to prove Section 3.1 from the main text. That is, that the objective values of the recommendations generated using joint knowledge gradient converge to the optimal values. The lions share of the work is done by the following proposition, which shows that the limiting covariance function $\overline{k}^{\infty}$ is pointwise almost surely constant. This is then used in Proposition A.15 to show that the limiting posterior mean $\overline{\mu}^{\infty}$ is pointwise almost surely equal to $\overline{f}$ up to a constant. The result is extended to show that the sample paths are almost surely equal up to a constant in Proposition A.16 and from this we deduce the convergence of the values of the recommendations the the supremum of $\overline{f}$.
@@
$\mathbb{P}\biggl{(}\overline{k}^{\infty}({\bm{x}},{\bm{g}};\,{\bm{x}}^{\prime},{\bm{g}}^{\prime})={\textnormal{c}}_{1}\biggr{)}=1.$
@@
By Lemma A.3, the posterior mean $\overline{\mu}^{n}({\bm{x}},{\bm{g}})=\mathbb{E}_{\mathbf{u}}[f({\bm{x}},{\bm{g}}({\mathbf{u}}),{\mathbf{u}})]$ attains its bounds, for each $n$. Thus, for each $n\in\mathbb{N}_{0}$, let ${\mathbf{x}}^{*n},{\mathbf{g}}^{*n}\in\operatorname*{arg\,max}_{{\mathbf{x}},{\mathbf{g}}}\overline{\mu}^{n}({\mathbf{x}},{\mathbf{g}})$ and for each $\tilde{{\bm{x}}}\in\tilde{{\mathbb{X}}}$, ${\bm{x}}^{\prime}\in{\mathbb{X}}$ and ${\bm{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}$ let
${\textnormal{w}}_{\tilde{{\bm{x}}}}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime})=\mathbb{E}\Bigl{[}\overline{f}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\overline{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\,\Big{|}\,\mathcal{F}^{n},{\textnormal{v}}_{\tilde{{\bm{x}}}}\Bigr{]}$
@@
${\textnormal{w}}_{\tilde{{\bm{x}}}}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime})$ $=\mathbb{E}\Bigl{[}\overline{f}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\overline{f}({\bm{x}}^{*n},{\bm{g}}^{*n})\,\Big{|}\,\mathcal{F}^{n},{\textnormal{v}}_{\tilde{{\bm{x}}}}\Bigr{]}$
@@
$=\overline{\mu}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\overline{\mu}^{n}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})+\frac{\mathbb{E}_{{\mathbf{u}}^{\prime}}[k^{n}(\tilde{{\mathbf{x}}}^{\prime},\tilde{{\bm{x}}})]-\mathbb{E}_{{\mathbf{u}}^{\prime}}[k^{n}(\tilde{{\mathbf{x}}}^{*n},\tilde{{\bm{x}}})]}{k^{n}(\tilde{{\bm{x}}},\tilde{{\bm{x}}})+\sigma^{2}}\bigl{(}{\textnormal{v}}_{\tilde{{\bm{x}}}}-\mu^{n}(\tilde{{\bm{x}}})\bigr{)}.$
@@
$\mathbb{E}[{\textnormal{w}}_{\tilde{{\bm{x}}}}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime})\,|\,\mathcal{F}^{n}]$ $=\overline{\mu}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\overline{\mu}^{n}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n}),$
@@
$\mathbb{P}({\textnormal{w}}_{\tilde{{\bm{x}}}}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime})>\delta\,|\,\mathcal{F}^{n})=1-\Phi\left(\frac{\sqrt{k^{n}(\tilde{{\bm{x}}},\tilde{{\bm{x}}})+\sigma^{2}}}{\bigl{|}\mathbb{E}_{{\mathbf{u}}^{\prime}}[k^{n}(\tilde{{\mathbf{x}}}^{\prime},\tilde{{\bm{x}}})]-\mathbb{E}_{{\mathbf{u}}^{\prime}}[k^{n}(\tilde{{\mathbf{x}}}^{*n},\tilde{{\bm{x}}})]\bigr{|}}\Bigl{[}\delta-\bigl{(}\overline{\mu}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\overline{\mu}^{n}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\bigr{)}\Bigr{]}\right).$
By Proposition A.12, $\overline{\mu}^{n}\to\overline{\mu}^{\infty}$ uniformly, almost surely as $n\to\infty$. Further, the same argument as in Lemma A.3 shows that $\overline{\mu}^{\infty}$ is bounded and attains its bounds. Therefore, there exists a negative random variable b (not depending on $n$) such that for sufficiently large $n$, $\overline{\mu}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\overline{\mu}^{n}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})>{\textnormal{b}}$. For example, take any ${\textnormal{b}}<\min\overline{\mu}^{\infty}-\max\overline{\mu}^{\infty}$. Then, almost surely,
@@
$\overline{k}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime};\,{\bm{x}},{\bm{g}})-\overline{k}^{n}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n};\,{\bm{x}},{\bm{g}})=\mathbb{E}_{{\mathbf{u}},{\mathbf{u}}^{\prime}}[k^{n}(\tilde{{\mathbf{x}}}^{\prime},\tilde{{\mathbf{x}}})-k^{n}(\tilde{{\mathbf{x}}}^{*n},\tilde{{\mathbf{x}}})]\to 0$
almost surely as $n\to\infty$. Further, from Proposition A.12, $\overline{k}^{n}\to\overline{k}^{\infty}$ uniformly, almost surely, so
@@
$\overline{k}^{\infty}({\bm{x}}^{\prime},{\bm{g}}^{\prime};\,{\bm{x}},{\bm{g}})-\overline{k}^{\infty}({\bm{x}},{\bm{g}};\,{\bm{x}},{\bm{g}})=\\ \Bigl{(}\overline{k}^{\infty}({\bm{x}}^{\prime},{\bm{g}}^{\prime};\,{\bm{x}},{\bm{g}})-\overline{k}^{\infty}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n};\,{\bm{x}},{\bm{g}})\Bigr{)}-\Bigl{(}\overline{k}^{\infty}({\bm{x}},{\bm{g}};\,{\bm{x}},{\bm{g}})-\overline{k}^{\infty}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n};\,{\bm{x}},{\bm{g}})\Bigr{)}\to 0\quad\text{a.s.}$
as $n\to\infty$. The left hand side is independent of $n$, so therefore must equal zero almost surely. Thus, by symmetry of $\overline{k}^{\infty}$, we conclude that for any ${\bm{x}}_{A},{\bm{x}}_{A}^{\prime},{\bm{x}}_{B},{\bm{x}}_{B}^{\prime}\in{\mathbb{X}}$ and ${\bm{g}}_{A},{\bm{g}}_{A}^{\prime},{\bm{g}}_{B},{\bm{g}}_{B}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}$,
$\overline{k}^{\infty}({\bm{x}}_{A},{\bm{g}}_{A};\,{\bm{x}}_{A}^{\prime},{\bm{g}}_{A}^{\prime})=\overline{k}^{\infty}({\bm{x}}_{B},{\bm{g}}_{B};\,{\bm{x}}_{B}^{\prime},{\bm{g}}_{B}^{\prime})$
almost surely. Taking ${\textnormal{c}}_{1}=\overline{k}^{\infty}({\bm{x}}_{A},{\bm{g}}_{A};\,{\bm{x}}_{A}^{\prime},{\bm{g}}_{A}^{\prime})$ for any choice of ${\bm{x}}_{A},{\bm{g}}_{A},{\bm{x}}_{A}^{\prime},{\bm{g}}_{A}^{\prime}$ completes the proof. ∎
@@
$\mathbb{P}\biggl{(}\overline{f}({\bm{x}},{\bm{g}})=\overline{\mu}^{\infty}({\bm{x}},{\bm{g}})+{\textnormal{c}}_{2}\biggr{)}=1.$

### v2

@@ A.3 Uniform convergence of the posterior mean and covariance functions
In this section, we establish the uniform convergence of the posterior mean and covariance functions of $f$ and $\bar{f}$. This will be an important ingredient in proving convergence of the values of the recommendations.
@@
The proof uses the theory of martingales in Banach spaces. The Gaussian process, $f$, is defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$ and can be viewed as a function $f:\Omega\times\tilde{{\mathbb{X}}}\to\mathbb{R},\,(\omega,\tilde{{\bm{x}}})\mapsto f(\tilde{{\bm{x}}})(\omega)$. The sigma-algebra $\mathcal{F}$ contains all the events generated by the finite-dimensional distributions of $f$ and, since sigma-algebras are closed under countable intersections, $\mathcal{F}$ also contains all events which can be described by the value of $f$ at countably many locations in $\tilde{{\mathbb{X}}}$. We are considering the version of $f$ with continuous sample paths and $\tilde{{\mathbb{X}}}$ is separable, and so $\mathcal{F}$ contains all events described by the sample paths of $f$. That is, we can view $f$ as a measurable function $f:\Omega\to\mathcal{C}(\tilde{{\mathbb{X}}}),\,\omega\mapsto(\tilde{{\bm{x}}}\mapsto f(\omega,\tilde{{\bm{x}}}))$ taking values in the space $\mathcal{C}(\tilde{{\mathbb{X}}})$ of continuous, real-valued functions on $\tilde{{\mathbb{X}}}$. In other words, $f$ is a random element in $\mathcal{C}(\tilde{{\mathbb{X}}})$, defined on $(\Omega,\mathcal{F},\mathbb{P})$.
@@
We now prove the analogous result of Proposition A.11 for the average performance objective, $\bar{f}$. For each $n\in\mathbb{N}_{0}\cup\{\infty\}$, ${\bm{x}},{\bm{x}}^{\prime}\in{\mathbb{X}}$ and ${\bm{g}},{\bm{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}$, define the posterior mean and covariance functions of $\bar{f}$ as
$\bar{\mu}^{n}({\bm{x}},{\bm{g}})$ $=\mathbb{E}[\bar{f}({\bm{x}},{\bm{g}})\,|\,\mathcal{F}^{n}],$ (30a)
$\bar{k}^{n}({\bm{x}},{\bm{g}};\,{\bm{x}}^{\prime},{\bm{g}}^{\prime})$ $=\mathrm{Cov}[\bar{f}({\bm{x}},{\bm{g}}),\bar{f}({\bm{x}}^{\prime},{\bm{g}}^{\prime})\,|\,\mathcal{F}^{n}].$ (30b)
Note that we do not include a statement of continuity of $\bar{\mu}^{\infty}$ and $\bar{k}^{\infty}$ since we have not defined a topology on the set ${\mathbb{X}}\times\{{\mathbb{U}}\to{\mathbb{Y}}\}$.
@@
For any choice of design points $\tilde{{\mathbf{x}}}^{1},\tilde{{\mathbf{x}}}^{2},\dots$, the sequences of stochastic processes $\bar{\mu}^{n}\to\bar{\mu}^{\infty}$ and $\bar{k}^{n}\to\bar{k}^{\infty}$ converge uniformly as $n\to\infty$, both almost surely and in $L^{p}$ for all $1\leq p<\infty$.
@@
$\sup_{\begin{subarray}{c}{\mathbf{x}}\in{\mathbb{X}}\\ {\mathbf{g}}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\bigl{|}\bar{\mu}^{n}({\mathbf{x}},{\mathbf{g}})-\bar{\mu}^{\infty}({\mathbf{x}},{\mathbf{g}})\bigr{|}\leq\sup_{\begin{subarray}{c}{\mathbf{x}}\in{\mathbb{X}}\\ {\mathbf{y}}\in{\mathbb{Y}}\\ {\mathbf{u}}\in{\mathbb{U}}\end{subarray}}\bigl{|}\mu^{n}({\mathbf{x}},{\mathbf{y}},{\mathbf{u}})-\mu^{\infty}({\mathbf{x}},{\mathbf{y}},{\mathbf{u}})\bigr{|}\to 0$
@@
$\sup_{\begin{subarray}{c}{\mathbf{x}},{\mathbf{x}}^{\prime}\in{\mathbb{X}}\\ {\mathbf{g}},{\mathbf{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\bigl{|}\bar{k}^{n}({\mathbf{x}},{\mathbf{g}};\,{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})-\bar{k}^{\infty}({\mathbf{x}},{\mathbf{g}};\,{\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})\bigr{|}\leq\sup_{\tilde{{\mathbf{x}}},\tilde{{\mathbf{x}}}^{\prime}\in\tilde{{\mathbb{X}}}}\bigl{|}k^{n}(\tilde{{\mathbf{x}}},\tilde{{\mathbf{x}}}^{\prime})-k^{\infty}(\tilde{{\mathbf{x}}},\tilde{{\mathbf{x}}}^{\prime})\bigr{|}\to 0$
@@
All moments of the supremum of $|\bar{f}|$ are finite. That is,
 $\forall\,1\leq p<\infty\qquad\mathbb{E}\Biggl{[}\biggl{(}\sup_{\begin{subarray}{c}{\mathbf{x}}\in{\mathbb{X}}\\ {\mathbf{g}}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}|\bar{f}({\mathbf{x}},{\mathbf{g}})|\biggr{)}^{p}\Biggr{]}<\infty.$
@@
$\mathbb{E}\Biggl{[}\biggl{(}\sup_{\begin{subarray}{c}{\mathbf{x}}\in{\mathbb{X}}\\ {\mathbf{g}}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}|\bar{f}({\mathbf{x}},{\mathbf{g}})|\biggr{)}^{p}\Biggr{]}=\mathbb{E}\Biggl{[}\biggl{(}\sup_{\begin{subarray}{c}{\mathbf{x}}\in{\mathbb{X}}\\ {\mathbf{g}}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\Bigl{|}\mathbb{E}_{\mathbf{u}}\bigl{[}f({\mathbf{x}},{\mathbf{g}}({\mathbf{u}}),{\mathbf{u}})\bigr{]}\Bigr{|}\biggr{)}^{p}\Biggr{]}\leq\mathbb{E}\biggl{[}\sup_{{\mathbf{x}},{\mathbf{y}},{\mathbf{u}}}|f({\mathbf{x}},{\mathbf{y}},{\mathbf{u}})|^{p}\biggr{]}<\infty.$
@@ A.4 Convergence to optimal recommendations
The main result in this section is to prove Section 3.1 from the main text. That is, that the objective values of the recommendations generated using joint knowledge gradient converge to the optimal values. The lions share of the work is done by the following proposition, which shows that the limiting covariance function $\bar{k}^{\infty}$ is pointwise almost surely constant. This is then used in Proposition A.15 to show that the limiting posterior mean $\bar{\mu}^{\infty}$ is pointwise almost surely equal to $\bar{f}$ up to a constant. The result is extended to show that the sample paths are almost surely equal up to a constant in Proposition A.16 and from this we deduce the convergence of the values of the recommendations the the supremum of $\bar{f}$.
@@
$\mathbb{P}\biggl{(}\bar{k}^{\infty}({\bm{x}},{\bm{g}};\,{\bm{x}}^{\prime},{\bm{g}}^{\prime})={\textnormal{c}}_{1}\biggr{)}=1.$
@@
By Lemma A.3, the posterior mean $\bar{\mu}^{n}({\bm{x}},{\bm{g}})=\mathbb{E}_{\mathbf{u}}[f({\bm{x}},{\bm{g}}({\mathbf{u}}),{\mathbf{u}})]$ attains its bounds, for each $n$. Thus, for each $n\in\mathbb{N}_{0}$, let ${\mathbf{x}}^{*n},{\mathbf{g}}^{*n}\in\operatorname*{arg\,max}_{{\mathbf{x}},{\mathbf{g}}}\bar{\mu}^{n}({\mathbf{x}},{\mathbf{g}})$ and for each $\tilde{{\bm{x}}}\in\tilde{{\mathbb{X}}}$, ${\bm{x}}^{\prime}\in{\mathbb{X}}$ and ${\bm{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}$ let
${\textnormal{w}}_{\tilde{{\bm{x}}}}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime})=\mathbb{E}\Bigl{[}\bar{f}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\bar{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\,\Big{|}\,\mathcal{F}^{n},{\textnormal{v}}_{\tilde{{\bm{x}}}}\Bigr{]}$
@@
${\textnormal{w}}_{\tilde{{\bm{x}}}}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime})$ $=\mathbb{E}\Bigl{[}\bar{f}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\bar{f}({\bm{x}}^{*n},{\bm{g}}^{*n})\,\Big{|}\,\mathcal{F}^{n},{\textnormal{v}}_{\tilde{{\bm{x}}}}\Bigr{]}$
@@
$=\bar{\mu}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\bar{\mu}^{n}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})+\frac{\mathbb{E}_{{\mathbf{u}}^{\prime}}[k^{n}(\tilde{{\mathbf{x}}}^{\prime},\tilde{{\bm{x}}})]-\mathbb{E}_{{\mathbf{u}}^{\prime}}[k^{n}(\tilde{{\mathbf{x}}}^{*n},\tilde{{\bm{x}}})]}{k^{n}(\tilde{{\bm{x}}},\tilde{{\bm{x}}})+\sigma^{2}}\bigl{(}{\textnormal{v}}_{\tilde{{\bm{x}}}}-\mu^{n}(\tilde{{\bm{x}}})\bigr{)}.$
@@
$\mathbb{E}[{\textnormal{w}}_{\tilde{{\bm{x}}}}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime})\,|\,\mathcal{F}^{n}]$ $=\bar{\mu}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\bar{\mu}^{n}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n}),$
@@
$\mathbb{P}({\textnormal{w}}_{\tilde{{\bm{x}}}}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime})>\delta\,|\,\mathcal{F}^{n})=1-\Phi\left(\frac{\sqrt{k^{n}(\tilde{{\bm{x}}},\tilde{{\bm{x}}})+\sigma^{2}}}{\bigl{|}\mathbb{E}_{{\mathbf{u}}^{\prime}}[k^{n}(\tilde{{\mathbf{x}}}^{\prime},\tilde{{\bm{x}}})]-\mathbb{E}_{{\mathbf{u}}^{\prime}}[k^{n}(\tilde{{\mathbf{x}}}^{*n},\tilde{{\bm{x}}})]\bigr{|}}\Bigl{[}\delta-\bigl{(}\bar{\mu}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\bar{\mu}^{n}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\bigr{)}\Bigr{]}\right).$
By Proposition A.12, $\bar{\mu}^{n}\to\bar{\mu}^{\infty}$ uniformly, almost surely as $n\to\infty$. Further, the same argument as in Lemma A.3 shows that $\bar{\mu}^{\infty}$ is bounded and attains its bounds. Therefore, there exists a negative random variable b (not depending on $n$) such that for sufficiently large $n$, $\bar{\mu}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\bar{\mu}^{n}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})>{\textnormal{b}}$. For example, take any ${\textnormal{b}}<\min\bar{\mu}^{\infty}-\max\bar{\mu}^{\infty}$. Then, almost surely,
@@
$\bar{k}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime};\,{\bm{x}},{\bm{g}})-\bar{k}^{n}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n};\,{\bm{x}},{\bm{g}})=\mathbb{E}_{{\mathbf{u}},{\mathbf{u}}^{\prime}}[k^{n}(\tilde{{\mathbf{x}}}^{\prime},\tilde{{\mathbf{x}}})-k^{n}(\tilde{{\mathbf{x}}}^{*n},\tilde{{\mathbf{x}}})]\to 0$
almost surely as $n\to\infty$. Further, from Proposition A.12, $\bar{k}^{n}\to\bar{k}^{\infty}$ uniformly, almost surely, so
@@
$\bar{k}^{\infty}({\bm{x}}^{\prime},{\bm{g}}^{\prime};\,{\bm{x}},{\bm{g}})-\bar{k}^{\infty}({\bm{x}},{\bm{g}};\,{\bm{x}},{\bm{g}})=\\ \Bigl{(}\bar{k}^{\infty}({\bm{x}}^{\prime},{\bm{g}}^{\prime};\,{\bm{x}},{\bm{g}})-\bar{k}^{\infty}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n};\,{\bm{x}},{\bm{g}})\Bigr{)}-\Bigl{(}\bar{k}^{\infty}({\bm{x}},{\bm{g}};\,{\bm{x}},{\bm{g}})-\bar{k}^{\infty}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n};\,{\bm{x}},{\bm{g}})\Bigr{)}\to 0\quad\text{a.s.}$
as $n\to\infty$. The left hand side is independent of $n$, so therefore must equal zero almost surely. Thus, by symmetry of $\bar{k}^{\infty}$, we conclude that for any ${\bm{x}}_{A},{\bm{x}}_{A}^{\prime},{\bm{x}}_{B},{\bm{x}}_{B}^{\prime}\in{\mathbb{X}}$ and ${\bm{g}}_{A},{\bm{g}}_{A}^{\prime},{\bm{g}}_{B},{\bm{g}}_{B}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}$,
$\bar{k}^{\infty}({\bm{x}}_{A},{\bm{g}}_{A};\,{\bm{x}}_{A}^{\prime},{\bm{g}}_{A}^{\prime})=\bar{k}^{\infty}({\bm{x}}_{B},{\bm{g}}_{B};\,{\bm{x}}_{B}^{\prime},{\bm{g}}_{B}^{\prime})$
almost surely. Taking ${\textnormal{c}}_{1}=\bar{k}^{\infty}({\bm{x}}_{A},{\bm{g}}_{A};\,{\bm{x}}_{A}^{\prime},{\bm{g}}_{A}^{\prime})$ for any choice of ${\bm{x}}_{A},{\bm{g}}_{A},{\bm{x}}_{A}^{\prime},{\bm{g}}_{A}^{\prime}$ completes the proof. ∎
@@
$\mathbb{P}\biggl{(}\bar{f}({\bm{x}},{\bm{g}})=\bar{\mu}^{\infty}({\bm{x}},{\bm{g}})+{\textnormal{c}}_{2}\biggr{)}=1.$